A point charge 50 microC is located in the XYZ plane at the point of position vector r=(2î+3j)meter. What is the electric field at the point of position vector r=(8î-5j)meter.

Grade:12

1 Answers

Arun
5 years ago
Charge is at (2,3)
for electric field at (8,-5) 
r = √(6×6+8×8) = 10
E = kq/r^2
= 9×10^9×50×10^-6÷100
= 450×10
=4500 Volt m